OpenStudy (anonymous):

Medal To whoever answers it!!! :) What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (anonymous):

slope of line =(7-3)/(2-(-2))=4/4=1 required equation of line is y-3=1(x-(-2)) y-3=x+4 y-x=4+3=7

OpenStudy (anonymous):

Thank you!!